A simple machine raises a load of 75kg through a distance of 0.6m. The effort applied to

the machine is 250N and moves through a distance of 18m. Taking g as 9.8m/s²,

determine, The mechanical advantage, The velocity ratio, & The efficiency of the machine.


1
Expert's answer
2021-10-14T18:41:42-0400

"M.A=L/E=75\times 9.8/250=2.94"

V.R= distance moved by effort"\/" distance moved by load="18m\/0.6m=30"

"efficiency=M.A/V.R \times 100\%=2.94/30\times 100\%=9.8\%"